# Self Model for Embodied Intelligence: Modeling Full-Body Human Musculoskeletal System and Locomotion Control with Hierarchical Low-Dimensional Representation

### 09 Dec 2023 v1

@@ V-B Baselines
We compared our algorithm with current state-of-the-art methods across three tasks: E2E, SAR, and SAR-ours. Our algorithm begins with the collection stage over a time step of length $M$ with muscle synergy, followed by optimization over a training stage of length $N$. One selecting of encoder-decoder architecture, inspired by SAR, we employed PCA[45] and ICA[46], where $u_{c}=\text{ICA}(\text{PCA}(u))$. In contrast, E2E is directly trained over a time step of $M+N$. SAR collects low-dimensional representations during a time step of $M$ and is trained over a time step of $N$. SAR-ours, on the other hand, extracts low-dimensional representations using muscle synergy during the time step of $M$ and then trains over a time step of length $N$.
@@ V-C Experimental Results
In our experiments, we utilized the SAC implementation from the stable-baselines3 framework[47]. For the Walking forward and Human-Exo interaction tasks, both the policy network and the value function network took dimensions of (512, 300). For Prosthetic walking, the dimensions were (400, 300), while all other parameters remained at their default settings. All observation values were normalized to a maximum value of 10. For the Walking forward task, the collection stage has a length ($M$) of 15M, while the training stage ($N$) is 20M. In the Human-Exo interaction and Prosthetic walking tasks, both $M$ and $N$ are set to 15M. The latent dimension $d$ is 50.
TABLE I: Comparing SOTA performance on the three tasks.
@@ V-D Analysis and Discussion
We demonstrate the visualization of our model’s motion performance in Fig. 2 and Fig. 5. We’ll also provide supplementary video. As shown in Fig. 5, in the human-exoskeleton interaction task, the algorithm is able to learn a policy that not only maintains upper body balance but also minimizes interaction forces between human and the exoskeleton.
@@ VI Conclusion
This work presents a comprehensive musculoskeletal model that encompasses the entire body. The model can be embedded into efficient simulation engines to provide a simulation framework with a generalized interface for reinforcement learning. We propose a new control algorithm designed for high-dimensional and nonlinear musculoskeletal models to generate biologically plausible human motion. This algorithm employs a hierarchical low-dimensional representation approach combined with deep reinforcement learning and imitation techniques. Remarkably surpassing baseline control algorithms, our proposed method demonstrates superior performance in full-body musculoskeletal model control tasks. Moreover, the derived control policy of the agents offers valuable insights into understanding muscle activation during locomotion.
This paper provides a full-body musculoskeletal model for the research community. The modeling and control of musculoskeletal systems has the potential to foster a deeper understanding of human motion control and human factors in human-machine interactions. As a human self-model for embodied intelligence, it could also serve as a test environment for the design of interactive robots and provide insights into humanoids.
@@
VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### 09 Mar 2024 v2

@@ V-B Baselines
We compared our algorithm with current state-of-the-art methods: E2E, SAR, and SAR-ours. Our algorithm begins with the collection stage over a time step of length $M$ with muscle synergy, followed by optimization over a training stage of length $N$. When selecting of encoder-decoder architecture, inspired by SAR, we employed PCA[45] and ICA[46], where $u_{c}=\text{ICA}(\text{PCA}(u))$. In contrast, E2E is directly trained over a time step of $M+N$. SAR collects low-dimensional representations during a time step of $M$ and is trained over a time step of $N$. SAR-ours, on the other hand, extracts low-dimensional representations using muscle synergy during the time step of $M$ and then trains over a time step of length $N$.
@@ V-C Experimental Results
In our experiments, we utilized the SAC implementation from the open-source RL library stable-baselines3[47]. For the Walking forward and Human-Exo interaction tasks, both the policy network and the value function network took dimensions of (512, 300). For Prosthetic walking, the dimensions were (400, 300), while all other parameters remained at their default settings. All observation values were normalized to a maximum value of 10. For the Walking forward task, the collection stage has a length ($M$) of 15M, while the training stage ($N$) is 20M. In the Human-Exo interaction and Prosthetic walking tasks, both $M$ and $N$ are set to 15M. The latent dimension $d$ is 50.
TABLE I: Comparing the maximum returns of algorithms on the three tasks.
@@ V-D Analysis and Discussion
We demonstrate the visualization of our model’s motion performance in Fig. 2 and Fig. 5. We also provide a supplementary video. As shown in Fig. 5, in the human-exoskeleton interaction task, the algorithm is able to learn a policy that not only maintains upper body balance but also minimizes interaction forces between human and the exoskeleton.
@@ VI Conclusion
This work introduces a musculoskeletal model including 90 rigid body segments, 206 joints, and 700 muscle-tendon units that encompasses the torso and limbs of human body. The model is adaptable to efficient open-source simulation engines, offering a simulation framework with a versatile interface for neuro-musculo-skeletal control simulation of human full-body movements. We introduce a novel control algorithm designed for high-dimensional and nonlinear musculoskeletal models to generate biologically plausible human motion. Our algorithm employs a hierarchical low-dimensional representation approach, integrating deep reinforcement learning and imitation techniques. Our proposed method exhibits superior performance in full-body musculoskeletal model control tasks, surpassing baseline algorithms significantly. Furthermore, the derived control policy of the agents provides valuable insights into the quantitative analysis of human motion control. Modeling and controlling musculoskeletal systems hold the potential to deepen our understanding of human motor intelligence and human factors in human-machine interactions. As a self-model of human for embodied intelligence, it could serve as a testing ground for the design of interactive robots and offer insights into humanoid behavior.